# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ I.1 Tiny-Imagenet Experiments
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03 95.91 $\pm$0.10 95.71 $\pm$0.03 97.61 $\pm$0.03
CIFAR-10.1 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02 92.97 $\pm$0.13 92.10 $\pm$0.33 93.45 $\pm$0.30
Gaussian Noise 64.64 68.87 $\pm$0.20 64.44 $\pm$0.11 70.04 $\pm$0.31 72.81 $\pm$0.09 72.73 $\pm$0.03 80.47 $\pm$0.26
Shot Noise 67.82 71.95 $\pm$0.06 66.81 $\pm$0.19 71.44 $\pm$0.16 74.45 $\pm$0.16 74.60 $\pm$0.03 82.64 $\pm$0.29
Impulse Noise 78.21 80.22 $\pm$0.19 76.46 $\pm$0.17 79.42 $\pm$0.15 81.36 $\pm$0.09 80.95 $\pm$0.15 87.80 $\pm$0.15
Defocus Blur 80.73 83.10 $\pm$0.03 79.01 $\pm$0.23 81.75 $\pm$0.19 83.20 $\pm$0.10 83.15 $\pm$0.18 90.33 $\pm$0.12
Glass Blur 50.29 57.12 $\pm$0.07 49.64 $\pm$0.23 58.13 $\pm$0.23 61.51 $\pm$0.07 62.35 $\pm$0.15 77.72 $\pm$0.31
Motion Blur 80.75 82.69 $\pm$0.11 78.85 $\pm$0.04 80.76 $\pm$0.12 82.60 $\pm$0.13 82.61 $\pm$0.12 88.93 $\pm$0.17
Zoom Blur 82.75 84.91 $\pm$0.08 82.32 $\pm$0.13 83.39 $\pm$0.05 85.76 $\pm$0.06 85.44 $\pm$0.13 91.69 $\pm$0.13
Snow 83.01 85.99 $\pm$0.11 82.69 $\pm$0.10 84.48 $\pm$0.07 84.91 $\pm$0.13 85.61 $\pm$0.15 91.56 $\pm$0.11
Frost 84.90 87.15 $\pm$0.12 84.63 $\pm$0.08 85.21 $\pm$0.06 87.17 $\pm$0.13 86.88 $\pm$0.04 91.43 $\pm$0.12
Fog 78.44 81.30 $\pm$0.07 77.56 $\pm$0.17 79.27 $\pm$0.07 81.80 $\pm$0.11 81.79 $\pm$0.09 90.82 $\pm$0.11
Brightness 91.67 93.07 $\pm$0.04 90.94 $\pm$0.04 91.87 $\pm$0.09 92.78 $\pm$0.05 92.59 $\pm$0.16 95.86 $\pm$0.00
Contrast 84.20 87.93 $\pm$0.04 82.88 $\pm$0.09 86.19 $\pm$0.06 87.54 $\pm$0.12 87.38 $\pm$0.02 94.53 $\pm$0.08
Elastic Transform 65.45 69.96 $\pm$0.12 64.81 $\pm$0.14 67.43 $\pm$0.24 71.19 $\pm$0.07 71.25 $\pm$0.09 82.16 $\pm$0.23
Pixelate 75.10 77.89 $\pm$0.05 72.92 $\pm$0.12 77.11 $\pm$0.10 77.88 $\pm$0.13 77.67 $\pm$0.16 87.13 $\pm$0.16
JPEG Compression 72.58 75.49 $\pm$0.07 71.18 $\pm$0.19 74.46 $\pm$0.11 75.88 $\pm$0.16 75.84 $\pm$0.18 82.60 $\pm$0.18
Mean 76.04 79.18 75.01 78.06 80.05 80.06 86.35
@@
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07 72.98 $\pm$0.07 72.85 $\pm$0.15 74.07 $\pm$0.20
Gaussian Noise 15.88 12.28 $\pm$0.20 15.43 $\pm$0.03 19.01 $\pm$0.24 34.23 $\pm$0.03 35.95 $\pm$0.27 38.80 $\pm$0.51
Shot Noise 17.49 15.07 $\pm$0.21 16.88 $\pm$0.07 20.27 $\pm$0.21 36.68 $\pm$0.10 37.96 $\pm$0.15 40.89 $\pm$0.10
Impulse Noise 21.43 13.13 $\pm$0.16 22.12 $\pm$0.15 17.66 $\pm$0.10 43.17 $\pm$0.35 44.62 $\pm$0.20 46.39 $\pm$0.17
Defocus Blur 40.10 50.35 $\pm$0.03 41.08 $\pm$0.22 49.86 $\pm$0.13 53.13 $\pm$0.12 53.80 $\pm$0.12 55.50 $\pm$0.24
Glass Blur 13.48 04.84 $\pm$0.14 18.43 $\pm$0.15 18.34 $\pm$0.31 32.53 $\pm$0.03 33.39 $\pm$0.11 37.89 $\pm$0.68
Motion Blur 39.82 49.85 $\pm$0.37 40.85 $\pm$0.26 50.00 $\pm$0.09 51.63 $\pm$0.06 52.72 $\pm$0.30 55.10 $\pm$0.28
Zoom Blur 45.45 54.76 $\pm$0.04 46.77 $\pm$0.06 54.13 $\pm$0.08 56.81 $\pm$0.11 57.51 $\pm$0.09 58.77 $\pm$0.33
Snow 42.77 52.38 $\pm$0.18 47.24 $\pm$0.18 52.80 $\pm$0.27 56.04 $\pm$0.06 56.73 $\pm$0.27 57.84 $\pm$0.13
Frost 45.39 51.66 $\pm$0.04 48.61 $\pm$0.14 49.56 $\pm$0.08 56.00 $\pm$0.11 56.48 $\pm$0.34 58.09 $\pm$0.04
Fog 38.98 50.74 $\pm$0.14 39.92 $\pm$0.16 49.92 $\pm$0.11 52.88 $\pm$0.33 53.83 $\pm$0.19 55.68 $\pm$0.38
Brightness 52.55 64.26 $\pm$0.09 55.83 $\pm$0.10 63.76 $\pm$0.13 65.58 $\pm$0.07 66.67 $\pm$0.19 67.76 $\pm$0.21
Contrast 33.32 48.69 $\pm$0.08 33.13 $\pm$0.16 47.86 $\pm$0.02 52.90 $\pm$0.06 55.06 $\pm$0.15 58.29 $\pm$0.51
Elastic Transform 24.39 33.56 $\pm$0.28 27.36 $\pm$0.10 32.93 $\pm$0.23 39.82 $\pm$0.21 40.37 $\pm$0.26 44.32 $\pm$0.18
Pixelate 21.89 36.20 $\pm$0.28 21.26 $\pm$0.10 39.49 $\pm$0.21 45.10 $\pm$0.06 47.02 $\pm$0.04 49.46 $\pm$0.17
JPEG Compression 27.21 30.80 $\pm$0.05 30.97 $\pm$0.10 35.56 $\pm$0.23 41.43 $\pm$0.18 42.13 $\pm$0.21 43.87 $\pm$0.23
Mean 32.01 37.90 33.73 40.08 47.86 48.95 51.24
@@
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08 79.33 $\pm$0.05 78.85 $\pm$0.19 81.38 $\pm$0.15
Gaussian Noise 30.55 36.93 $\pm$0.03 36.10 $\pm$0.11 41.46 $\pm$0.15 43.99 $\pm$0.13 44.13 $\pm$0.11 52.26 $\pm$0.18
Shot Noise 34.58 40.96 $\pm$0.16 38.23 $\pm$0.13 44.27 $\pm$0.09 46.28 $\pm$0.22 46.63 $\pm$0.17 54.59 $\pm$0.26
Impulse Noise 44.89 49.09 $\pm$0.14 49.69 $\pm$0.21 51.44 $\pm$0.23 56.15 $\pm$0.04 56.26 $\pm$0.22 63.02 $\pm$0.12
Defocus Blur 48.88 55.23 $\pm$0.07 50.43 $\pm$0.19 56.55 $\pm$0.22 57.46 $\pm$0.01 57.66 $\pm$0.42 65.20 $\pm$0.17
Glass Blur 23.46 27.02 $\pm$0.23 24.35 $\pm$0.22 30.47 $\pm$0.14 32.54 $\pm$0.12 33.54 $\pm$0.16 48.19 $\pm$0.17
Motion Blur 50.83 56.03 $\pm$0.20 51.94 $\pm$0.04 56.98 $\pm$0.18 58.22 $\pm$0.10 57.81 $\pm$0.05 65.09 $\pm$0.17
Zoom Blur 56.02 61.19 $\pm$0.10 56.96 $\pm$0.16 62.56 $\pm$0.04 62.94 $\pm$0.02 62.74 $\pm$0.06 68.92 $\pm$0.15
Snow 49.03 55.60 $\pm$0.09 54.89 $\pm$0.11 58.81 $\pm$0.11 60.68 $\pm$0.06 61.04 $\pm$0.27 66.81 $\pm$0.15
Frost 53.27 58.21 $\pm$0.15 58.15 $\pm$0.33 60.38 $\pm$0.23 62.34 $\pm$0.14 62.76 $\pm$0.22 67.71 $\pm$0.29
Fog 48.51 53.37 $\pm$0.25 49.26 $\pm$0.13 54.38 $\pm$0.04 54.71 $\pm$0.31 54.70 $\pm$0.13 65.54 $\pm$0.23
Brightness 60.53 67.34 $\pm$0.17 66.60 $\pm$0.10 69.63 $\pm$0.14 71.52 $\pm$0.07 71.60 $\pm$0.09 76.25 $\pm$0.15
Contrast 50.24 59.91 $\pm$0.13 53.64 $\pm$0.24 63.39 $\pm$0.13 62.77 $\pm$0.22 63.95 $\pm$0.04 72.34 $\pm$0.27
Elastic Transform 35.07 38.49 $\pm$0.12 35.72 $\pm$0.09 39.57 $\pm$0.39 41.28 $\pm$0.25 41.27 $\pm$0.15 52.72 $\pm$0.24
Pixelate 43.86 48.37 $\pm$0.17 44.32 $\pm$0.10 50.45 $\pm$0.16 51.15 $\pm$0.15 51.22 $\pm$0.13 59.55 $\pm$0.06
JPEG Compression 39.11 44.42 $\pm$0.09 43.44 $\pm$0.11 47.45 $\pm$0.14 49.40 $\pm$0.17 49.78 $\pm$0.18 54.98 $\pm$0.20
Mean 44.59 50.14 47.58 52.52 54.10 54.34 69.48
@@
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.12 63.59 63.87 $\pm$0.09 64.73 $\pm$0.11 66.05 $\pm$0.03 66.34 $\pm$0.07
Gaussian Noise 4.80 13.44 $\pm$0.21 10.53 $\pm$0.25 19.25 $\pm$0.18 16.78 $\pm$0.09 24.88 $\pm$0.28
Shot Noise 6.44 17.46 $\pm$0.20 12.90 $\pm$0.45 21.80 $\pm$0.11 19.27 $\pm$0.13 27.99 $\pm$0.32
Impulse Noise 4.36 9.09 $\pm$0.14 6.57 $\pm$0.20 13.46 $\pm$0.01 10.42 $\pm$0.07 20.69 $\pm$0.18
Defocus Blur 22.23 32.33 $\pm$0.09 29.16 $\pm$0.32 33.49 $\pm$0.29 $3$$\pm$0.03 37.27 $\pm$0.16
Glass Blur 7.50 12.47 $\pm$0.14 10.88 $\pm$0.18 18.33 $\pm$0.09 16.26 $\pm$0.08 22.72 $\pm$0.14
Motion Blur 33.46 44.17 $\pm$0.06 39.91 $\pm$0.43 45.17 $\pm$0.11 44.21 $\pm$0.10 48.75 $\pm$0.37
Zoom Blur 29.86 39.24 $\pm$0.26 36.62 $\pm$0.11 40.53 $\pm$0.19 40.26 $\pm$0.21 45.16 $\pm$0.08
Snow 31.64 38.91 $\pm$0.29 38.66 $\pm$0.19 42.71 $\pm$0.11 42.31 $\pm$0.18 45.74 $\pm$0.15
Frost 33.26 42.04 $\pm$0.15 40.95 $\pm$0.20 44.28 $\pm$0.14 44.83 $\pm$0.13 48.09 $\pm$0.05
Fog 24.26 31.06 $\pm$0.18 29.22 $\pm$0.24 34.78 $\pm$0.11 34.08 $\pm$0.13 41.21 $\pm$0.19
Brightness 40.48 49.26 $\pm$0.12 47.78 $\pm$0.44 51.61 $\pm$0.39 52.16 $\pm$0.08 55.75 $\pm$0.24
Contrast 1.72 3.06 $\pm$0.09 3.58 $\pm$0.07 7.24 $\pm$0.09 5.43 $\pm$0.08 13.06 $\pm$0.18
Elastic Transform 23.85 34.57 $\pm$0.05 32.11 $\pm$0.10 35.16 $\pm$0.05 35.11 $\pm$0.05 41.69 $\pm$0.12
Pixelate 22.35 37.73 $\pm$0.18 29.02 $\pm$0.22 41.11 $\pm$0.18 37.47 $\pm$0.21 45.83 $\pm$0.17
JPEG Compression 27.65 41.92 $\pm$0.45 36.51 $\pm$0.12 44.58 $\pm$0.04 42.93 $\pm$0.08 46.55 $\pm$0.26
Mean 20.92 29.78 26.96 32.90 31.66 37.69
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 71.34 74.09 $\pm$0.12 74.51 $\pm$0.20 74.57 $\pm$0.15 75.04 $\pm$0.21 75.01 $\pm$0.09
Gaussian Noise 17.26 26.60 $\pm$0.16 23.22 $\pm$0.12 29.27 $\pm$0.03 27.98 $\pm$0.14 38.14 $\pm$0.09
Shot Noise 21.57 28.93 $\pm$0.27 27.39 $\pm$0.21 33.01 $\pm$0.18 31.47 $\pm$0.04 41.02 $\pm$0.13
Impulse Noise 17.32 $2$$\pm$0.29 19.79 $\pm$0.18 27.93 $\pm$0.09 25.38 $\pm$0.15 35.48 $\pm$0.10
Defocus Blur 35.63 38.66 $\pm$0.19 39.44 $\pm$0.27 39.99 $\pm$0.15 40.55 $\pm$0.04 46.03 $\pm$0.17
Glass Blur 13.62 10.98 $\pm$0.13 15.72 $\pm$0.18 17.09 $\pm$0.15 17.44 $\pm$0.16 27.93 $\pm$0.18
Motion Blur 50.11 52.53 $\pm$0.01 53.91 $\pm$0.32 53.63 $\pm$0.28 54.46 $\pm$0.16 57.78 $\pm$0.06
Zoom Blur 43.65 46.02 $\pm$0.15 48.82 $\pm$0.11 47.39 $\pm$0.04 48.28 $\pm$0.09 52.67 $\pm$0.09
Snow 48.04 49.13 $\pm$0.33 51.50 $\pm$0.27 51.97 $\pm$0.14 52.60 $\pm$0.04 56.85 $\pm$0.13
Frost 49.36 50.46 $\pm$0.09 53.89 $\pm$0.22 51.49 $\pm$0.11 53.77 $\pm$0.16 57.56 $\pm$0.35
Fog 37.58 38.63 $\pm$0.15 40.43 $\pm$0.11 42.35 $\pm$0.04 42.57 $\pm$0.09 50.49 $\pm$0.09
Brightness 55.85 59.25 $\pm$0.04 61.07 $\pm$0.09 60.69 $\pm$0.12 62.01 $\pm$0.04 65.28 $\pm$0.31
Contrast 5.78 8.66 $\pm$0.14 7.48 $\pm$0.18 11.25 $\pm$0.08 9.21 $\pm$0.15 21.57 $\pm$0.13
Elastic Transform 34.15 36.81 $\pm$0.19 40.65 $\pm$0.27 38.44 $\pm$0.22 39.85 $\pm$0.08 47.86 $\pm$0.17
Pixelate 45.90 50.77 $\pm$0.18 52.28 $\pm$0.31 52.22 $\pm$0.06 51.83 $\pm$0.04 58.23 $\pm$0.29
JPEG Compression 50.13 51.88 $\pm$0.13 52.38 $\pm$0.24 52.46 $\pm$0.19 54.19 $\pm$0.16 56.82 $\pm$0.24
Mean 34.98 40.28 41.07 42.98 43.28 50.36
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-Imagenet 58.29 57.72 $\pm$0.14 58.90 $\pm$0.15 59.85 $\pm$0.01 61.35 $\pm$0.17 63.69 $\pm$0.15
Gaussian Noise 7.08 8.01 $\pm$0.05 9.29 $\pm$0.03 14.44 $\pm$0.25 13.02 $\pm$0.07 21.40 $\pm$0.37
Shot Noise 9.41 10.04 $\pm$0.07 11.70 $\pm$0.07 17.44 $\pm$0.12 15.94 $\pm$0.07 24.90 $\pm$0.17
Impulse Noise 3.44 4.18 $\pm$0.02 4.85 $\pm$0.01 10.37 $\pm$0.16 6.90 $\pm$0.01 17.34 $\pm$0.14
Defocus Blur 21.71 24.53 $\pm$0.05 27.56 $\pm$0.10 31.46 $\pm$0.26 29.91 $\pm$0.13 35.39 $\pm$0.16
Glass Blur 9.12 10.09 $\pm$0.06 11.03 $\pm$0.10 15.84 $\pm$0.09 14.01 $\pm$0.10 21.16 $\pm$0.19
Motion Blur 34.52 36.94 $\pm$0.07 38.97 $\pm$0.05 41.34 $\pm$0.20 41.26 $\pm$0.04 46.26 $\pm$0.37
Zoom Blur 27.44 29.48 $\pm$0.06 34.29 $\pm$0.07 35.06 $\pm$0.15 33.96 $\pm$0.07 40.93 $\pm$0.05
Snow 32.51 32.20 $\pm$0.03 34.45 $\pm$0.10 36.79 $\pm$0.08 37.76 $\pm$0.09 42.32 $\pm$0.14
Frost 36.33 35.72 $\pm$0.02 37.13 $\pm$0.05 38.37 $\pm$0.13 39.65 $\pm$0.03 44.97 $\pm$0.17
Fog 25.94 27.46 $\pm$0.04 28.89 $\pm$0.08 33.51 $\pm$0.19 32.13 $\pm$0.08 38.60 $\pm$0.32
Brightness 40.15 39.79 $\pm$0.08 43.31 $\pm$0.10 46.52 $\pm$0.06 46.93 $\pm$0.13 53.10 $\pm$0.05
Contrast 1.81 2.24 $\pm$0.04 3.15 $\pm$0.04 6.07 $\pm$0.12 3.53 $\pm$0.02 11.88 $\pm$0.23
Elastic Transform 30.40 31.92 $\pm$0.05 33.88 $\pm$0.14 33.74 $\pm$0.11 35.01 $\pm$0.13 40.73 $\pm$0.16
Pixelate 22.78 24.79 $\pm$0.03 27.70 $\pm$0.12 34.84 $\pm$0.09 31.55 $\pm$0.08 41.84 $\pm$0.09
JPEG Compression 29.59 30.93 $\pm$0.13 33.60 $\pm$0.15 37.29 $\pm$0.14 36.46 $\pm$0.11 42.86 $\pm$0.04
Mean 22.14 23.22 25.32 28.87 27.87 34.91
[…]

### v3

@@ D Comparison to OT-VP
Second Setting : Multi Source
@@
Our setting: No specific source
SAT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
E Discussion on the Two Hyperparameters of WATT and Motivation for SAT
WATT [osowiechi2024watt] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process). While these hyperparameters play an essential role in improving WATT’s adaptation capabilities, they introduce significant computational and scalability challenges, particularly as the number of classes grows. The iterative nature of $L$ requires multiple forward and backward passes for every template, leading to substantial runtime overhead, especially for datasets with many classes like Tiny-ImageNet. Additionally, higher $L$ values, while effective for complex corruptions, risk overfitting to noisy pseudo-labels, while lower $L$ values may lead to under-adaptation. Similarly, $M$ stabilizes adaptation through repeated weight averaging, but its effectiveness diminishes beyond a certain point, with larger values providing negligible performance improvements while exponentially increasing runtime. This dependence on repeated updates for each class and template becomes especially prohibitive for datasets with a large number of classes, as demonstrated in Figure 4, where runtime increases from CIFAR-10 (10 classes) to Tiny-ImageNet (200 classes).
These limitations motivate our proposed approach, SAT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$. SAT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging. This not only drastically reduces computational overhead but also ensures scalable, robust adaptation across diverse corruption levels and datasets, offering a practical alternative to WATT’s hyperparameter-dependent framework.
F Differences wrt PLOT [chenplot]
Note that despite the apparent similarities, our work presents fundamental differences w.r.t. Prompt Learning with Optimal Transport (PLOT) [chenplot], which prevents from adding the later in the empirical evaluation. a) Target task. PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available. Thus PLOT cannot be compared in the TTA setting. b) Adaptation. Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category). c) Use of OT. PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities. In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.